# Iwin Transformer: Hierarchical Vision Transformer using Interleaved Windows

### v1

@@ IV-F Ablation Study
We explored the impact of varying kernel sizes for DWConv across different stages. Our final configuration, which utilizes a fixed kernel size of {3, 3, 3, None}, yields a accuracy of 82.0% with a throughput of 736 img/s. While a {5, 5, 5, None} kernel size achieves a slightly higher accuracy of 82.2%, the smaller {3, 3, 3, None} kernels offer a more favorable trade-off between performance and computational efficiency (higher throughput). Additionally, using different kernel sizes at different stages to satisfy $KM\geq\max(H,W)$ did not yield the best results. This aligns with observations in [37] that balanced network outperform theoretically optimal but imbalanced configurations.
@@
TABLE VIII: Ablation studies for Iwin-T on the COCO with the Mask R-CNN 1$\times$ schedule. ”lr” denotes the initial learning rate. ”rel. pos.” indicates the relative position encoding.
@@ V-A Migration to Large Language Models
The Iwin Transformer’s position-embedding-free design principle offers promising opportunities for application in Large Language Models (LLMs). Currently, LLMs heavily depend on position embeddings to preserve sequence order information. By integrating interleaved window attention with depthwise separable convolution, it may be possible to achieve more natural length generalization. This approach relies on structural rather than parametric position information, facilitating easier length extrapolation. As shown in Figure 10, the computation is divided into two components: 1D causal depthwise separable convolution and 1D interleaved window causal attention. Both ensure tokens relate only to preceding tokens, giving Iwin 1D Attention causality. Furthermore, we can replace the depthwise separable convolution with normal window causal attention by setting two window sizes $M_{1}$ and $M_{2}$ equal, with $M_{1}M_{2}=N$, resulting in window size $\sqrt{N}$. This reduces complexity for sequence length $N$ from $N^{2}$ to $N$.

### v2

@@ IV-F Ablation Study
We explored the impact of varying kernel sizes for DWConv across different stages. Our final configuration, which utilizes a fixed kernel size of {3, 3, 3, None}, yields a accuracy of 82.0% with a throughput of 736 img/s. While a {5, 5, 5, None} kernel size achieves a slightly higher accuracy of 82.2%, the smaller {3, 3, 3, None} kernels offer a more favorable trade-off between performance and computational efficiency (higher throughput). Additionally, using different kernel sizes at different stages to satisfy $KM\geq\max(H,W)$ did not yield the best results. This aligns with observations in [tan2019efficientnet] that balanced network outperform theoretically optimal but imbalanced configurations.
@@
[…]
@@ V-A Migration to Large Language Models
The Iwin Transformer’s position-embedding-free design principle offers promising opportunities for application in Large Language Models (LLMs). Currently, LLMs heavily depend on position embeddings to preserve sequence order information. By integrating interleaved window attention with depthwise separable convolution, it may be possible to achieve more natural length generalization. This approach relies on structural rather than parametric position information, facilitating easier length extrapolation. As shown in Figure 12, the computation is divided into two components: 1D causal depthwise separable convolution and 1D interleaved window causal attention. Both ensure tokens relate only to preceding tokens, giving Iwin 1D Attention causality. Furthermore, we can replace the depthwise convolution with normal window causal attention by setting two window sizes $M_{1}$ and $M_{2}$ equal, with $M_{1}M_{2}=N$, resulting in window size $\sqrt{N}$. This reduces complexity for sequence length $N$ from $N^{2}$ to $N$.
@@ VI Conclusion
[Uncaptioned image] Simin Huo received his B.S. degree from Nanjing University of Science and Technology, China, Nanjing, in 2018, and his M.S. degree from Bauman Moscow State Technical University, Russia, Moscow, in 2021. He is currently a Ph.D. student with the Department of Automatics, Shanghai Jiao Tong University. His main research interests include computer vision and deep learning , with a focus on efficient vision transformer and generation models.
[Uncaptioned image] Ning Li (Member, IEEE) received the B.S. and M.S. degrees from Qingdao University of Science and Technology, Qingdao, China, in 1996 and 1999, respectively, and the Ph.D. degree from Shanghai Jiao Tong University, Shanghai, China, in 2002. She is currently a Professor with the Department of Automation, Shanghai Jiao Tong University. Her research interests include modeling and control of complex systems, artificial intelligence, and big data analysis.